# arXiv:2601.10941v2[math.AP] 01 Mar 2026

### v2

@@ Abstract.
We consider abstract nonlinear equations of the form $Au=\lambda u+I^{\prime}(u)$, where $A$ is a self-adjoint operator with compact resolvent on a Hilbert space $H$, $\lambda\in\mathbb{R}$ is a parameter, and $u\mapsto I^{\prime}(u)$ is a superlinear term of variational nature. In this abstract setting, we develop a new approach to existence and multiplicity of solutions with prescribed norm in $H$. We then consider various applications of this approach. First, we obtain, under fairly general assumptions including the mass-supercritical case, the existence of infinitely many solutions to a class of nonlinear (time-independent) Schrödinger equations on a compact graph $\mathcal{G}$ with prescribed (arbitrarily large) mass. Moreover, we derive a similar existence and multiplicity result for a biharmonic semilinear equation in the $2$-torus. For a larger class of second order and higher order equations in a bounded domain with Dirichlet boundary conditions, we also show the existence of multiple solutions with prescribed small mass.
The solutions we obtain are detected as ground states of Nehari-Pankov type for the associated $\lambda$-dependent action functional, where $\lambda$ varies in a spectral gap between sufficiently large eigenvalues of $A$. The key observation in this abstract framework is the fact that the $H$-norms of these $\lambda$-dependent solution families form connected sets. To estimate the size of these connected sets in specific settings, we use Weyl type estimates for the length of spectral gaps, variational characterizations of eigenvalues, bounds for associated eigenfunctions and a classical bound from analytic number theory.
@@ 1. Introduction
The assumption (f1) is a natural superlinearity condition. The condition (f2) can be seen a mild form of approximate homogeneity. Since, by (f1), the function $F$ is increasing on $\mathopen{[}0,+\infty\mathclose{)}$ and decreasing on $(-\infty,0]$, one can always replace $\kappa_{0}$ by a bigger positive constant in (4).
@@
Very recently, in [16, Theorem 1.3], Jeanjean and Song proved a multiplicity result for compact graphs $\mathcal{G}$ and small masses with continuity-Kirchhoff conditions. More precisely, by combining flow invariance techniques with delicate estimates, they show that for every $n\in\mathbb{Z}^{>0}$ there exists $\mu_{n}>0$ with the property that, for every $\mu\in(0,\mu_{n})$, the NLS equation $u^{\prime\prime}=\lambda u+|u|^{p-2}u$ admits at least $n$ pairs $\pm u$ of sign changing solutions with fixed mass $\mu$. Using our new “frequency to masses” tool (Theorem 1.8 below), we are able to improve this multiplicity result to the existence of infinitely many solutions, for any mass. Moreover, we can extend this multiplicity result beyond the class of odd nonlinearities to a setting where, in particular, Lusternik-Schnirelman theory does not apply. Let us also point out that we can deal with the most general self-adjoint vertex conditions, also generalizing the previous results. In the case of continuity-Kirchhoff conditions, the first eigenfunction of $A$ is positive in $\mathcal{G}$ (see e.g. [18, Theorem 4.12]), so that we also obtain the existence of sign-changing solutions. Finally, as already pointed out, we do not need to make any case distinction depending on whether $p<6$ or $6<p$.
@@
We emphasize that this result is new even in the classical second order case $m=1$, and it does not require oddness of the nonlinearity $f$ in $t$. In the special case of homogeneous nonlinearities of the form $f(x,t)=|t|^{p-2}t$, Theorem 1.3 is the analogue of the recent multiplicity result of Jeanjean and Song in [16, Theorem 1.3] for the case of compact metric graphs. Our method is completely different from the approach in [16], which, according to remarks in [16, Introduction], also yields multiplicity results for NLS-equations on bounded domains in $\mathbb{R}^{N}$.
@@
(i) In the mass supercritical regime $p>6$ and for a given positive mass $\mu>0$, even the mere existence of one normalized solution of (8) is new. If the weight function $r$ is constant, then the symmetries of the torus can be used to produce multiple solutions with large mass via a reflection procedure (see [21, Theorem 1.13 and Remark 1.1] for related results in bounded domains). This is not possible when $r$ does not have any symmetries.
@@
with strict inequality if $v=0$ and $s\neq 0$, see [25, Lemma 38]. This is very useful in applications.
@@
We emphazise that the connectedness of the set $M_{n}$ stated in Theorem 1.8 does not follow from a standard continuity argument, since we cannot expect connectedness of the set of associated ground state solutions in this general variational framework. Instead, the connectedness of $M_{n}$ follows by an inspection of the special saddle point nature of the Nehari-Pankov sets $\mathcal{N}\!{\mathcal{P}}_{\lambda}$, while it is not a sole consequence of the continuity of ground state levels $c_{\lambda}$.
@@
Let $k$ be a positive integer, and let $I_{1},\dots,I_{k}$ be disjoint spectral gaps [^3] We call an interval $I\subset\mathbb{R}$ a spectral gap of AA if $I=(\lambda_{n},\lambda_{n+1})$ for some $n\in\mathbb{Z}^{>0}$. of $A$. Then, there exists $\mu_{k}>0$ such that, for all $\mu\in(0,\mu_{k})$, there exist $\Lambda_{1},\dotsc,\Lambda_{k}\in\mathbb{R}$ and $u_{1},\dotsc,u_{k}\in E$ with
@@
In particular, we are able to prove Theorem 1.1 by exploiting the presence of large spectral gaps $I$ for Laplacians on compact metric graphs (a consequence of Weyl’s law, which holds for all self-adjoint realizations of the Laplacian on graphs as shown in [6]). Classical spectral theory then shows that for parameters $\lambda$ at the bottom of a large spectral gap, solutions of (3) have a large $L^{\infty}$-norm. We then conclude using an ODE argument providing a suitable $L^{\infty}$ to $L^{2}$” estimate.
@@
The paper is organized as follows. In Section 2, we further develop the abstract framework of the indefinite variational problem associated with (1) and prove Theorem 1.7. In Section 3, we then study the $\lambda$-dependence of the Nehari-Pankov solutions obtained in Theorem 1.7, and we complete the proof of Theorem 1.8. In the end of Section 3, we then provide useful estimates for the action and mass of Nehari-Pankov solutions under additional assumptions on the functional $I$ and the operator $A$. Section 5 is devoted to the nonlinear Schrödinger equation on compact metric graphs and to the proof of Theorem 1.1. Finally, in Section 6, we study the equation (8) on the flat $2$-torus, and we complete the proof of Theorem 1.4.
@@ 2. Existence of Nehari-Pankov solutions for fixed $\lambda$
Next, arguing as in [25, Chapter 4], we show that $\mathcal{N}\!{\mathcal{P}}_{\lambda}$ contains all nonzero critical points of $\mathcal{J}_{\lambda}$. Let $u\in E\setminus\{0\}$ be a critical point of $\mathcal{J}_{\lambda}$. Then, according to (20),
@@
(21) $\sup_{v\in E_{n}}\mathcal{J}_{\lambda}(v)\leq\sup_{v\in E_{n}}\frac{q_{\lambda}(v)}{2}\leq 0.\textasciidieresis$
@@
$w_{k}:=\frac{u_{k}}{t_{k}}\to w\qquad\text{in $E$ as $n\to\infty$,}$
@@
$\inf_{\mathcal{N}\!{\mathcal{P}}_{\lambda}}\mathcal{J}_{\lambda}=\inf_{\tilde{u}\in E\setminus E_{n}}\sup_{F(\tilde{u})^{+}}\mathcal{J}_{\lambda}.$
@@
Without loss of generality, we may assume that $w,w_{k}\in E_{n}^{\perp}$, and that $\|w\|=1$, $\|w_{k}\|_{E}=1$ for all $k$. We write
@@
$\mathcal{J}_{\lambda}(su+v)=\lim_{k\to\infty}\mathcal{J}_{\lambda}(su_{k}+v)\leq\lim_{k\to\infty}\mathcal{J}_{\lambda}(u_{k})=\mathcal{J}_{\lambda}(u).$
Consequently, $u$ maximizes $\mathcal{J}_{\lambda}$ on $F^{+}(u)=F^{+}(w)$, which implies that
@@ 3.1. Norms of Nehari-Pankov solutions
We first show (33). Let $\delta>0$ and $u\in Q_{\lambda}$. Then we have
$c_{\lambda}=\mathcal{J}_{\lambda}(u)\geq\mathcal{J}_{\lambda}(w)\qquad\text{for all $w\in\overline{F_{n}(u)^{+}}$}$
by Proposition 2.2. We claim that there exists $\varepsilon_{0}>0$ and $\kappa>0$ with
(35) $\mathcal{J}_{\lambda-\varepsilon}(w)\leq c_{\lambda}-\kappa\qquad\text{for all $\varepsilon\in(0,\varepsilon_{0})$ and $w\in F^{+}(u)\setminus B_{\delta}(u)$,}$
@@
$\lim_{n\to\infty}J_{\lambda-\varepsilon_{n}}(w_{n})\geq c_{\lambda}$
for a sequence $\varepsilon_{n}\to 0^{+}$ and a sequence $(w_{k})_{k}$ in $F^{+}(u)\setminus B_{\delta}(u)$. By Lemma 2.1 and the positivity of $c_{\lambda}$, the sequence $(w_{k})_{k}$ is bounded in $E$. Hence, since $F^{+}(u)\subset F_{n}(u)$ and the space $F_{n}(u)$ is finite dimensional, we may pass to a subsequence such that $w_{n}\to w$ in $E$ for some $w\in\overline{F^{+}(u)}\setminus B_{\delta}(u)$, which implies that
$\mathcal{J}_{\lambda}(w)=\lim_{n\to\infty}\mathcal{J}_{\lambda-\varepsilon_{n}}(w_{n})\geq c_{\lambda}>0.$
Consequently, $w\in F^{+}(u)\setminus B_{\delta}(u)$ by (21), but then the uniqueness stated in Proposition 2.2 gives that $u=w\in F^{+}(u)\setminus B_{\delta}(u)$, a contradiction. Hence (35) is true.
By the monotonicity property stated in Remark 3.1, we now observe that, for $\varepsilon\in(0,\varepsilon_{0})$,
$c_{\lambda}$ $\leq c_{\lambda-\varepsilon}\leq\sup_{F^{+}(u)}\mathcal{J}_{\lambda-\varepsilon}=\sup_{F^{+}(u)\cap B_{\delta}(u)}\mathcal{J}_{\lambda-\varepsilon}$
$\leq\sup_{F^{+}(u)}\mathcal{J}_{\lambda}+\frac{\varepsilon}{2}\sup_{B_{\delta}(u)}\|\cdot\|_{H}^{2}\leq c_{\lambda}+\frac{\varepsilon}{2}\Bigl(\|u\|_{H}+\delta\Bigr)^{2}$
@@
From (33) and the fact that the map $\lambda\mapsto c_{\lambda}$ is decreasing, it immediately follows that this map is left continuous. Next, to prove its right continuity, let $\lambda\in(\lambda_{n},\lambda_{n+1})$, and let $(\lambda_{n})_{n}$ be a sequence in $(\lambda,\lambda_{n+1})$ which converges to $\lambda$. Since the map $\lambda\mapsto c_{\lambda}$ is decreasing, it follows that
$c_{\lambda}\geq\limsup_{n\to\infty}c_{\lambda_{n}}.$
Let, moreover, $w_{n}\in E_{n}^{\perp}$ be chosen with $\|w_{n}\|_{E}=1$ and $\sup\limits_{F_{n}(w_{n})^{+}}\mathcal{J}_{\lambda_{n}}=c_{\lambda_{n}}$. From Lemma 2.4, it then follows that
$w_{n}\rightharpoonup w\in E\qquad\text{after passing to subsequences}$
@@
$c_{\lambda}\leq\sup_{F_{n}(w)^{+}}\mathcal{J}_{\lambda}\leq\liminf_{n\to\infty}\sup\limits_{F_{n}(w_{n})^{+}}\mathcal{J}_{\lambda_{n}}=\liminf_{n\to\infty}c_{\lambda_{n}}.$
This shows that $\lim\limits_{n\to\infty}c_{\lambda_{n}}=c_{\lambda}$, and we thus obtain the full continuity of the map $\lambda\mapsto c_{\lambda}$.
Finally, to show (34), we again let $u\in Q_{\lambda}$ and $\delta\in\mathopen{(}0,\|u\|_{H}\mathclose{)}$. Using the continuity of the map $\lambda\mapsto c_{\lambda}$ and arguing similarly as before, we see that there exist $\varepsilon_{0}>0$ and $\kappa>0$ with
$c_{\lambda+\varepsilon}>c_{\lambda}-\kappa\qquad\text{and}\qquad\mathcal{J}_{\lambda+\varepsilon}(w)\leq c_{\lambda}-\kappa\qquad\text{for $\varepsilon\in(0,\varepsilon_{0})$ and $w\in F^{+}(u)\setminus B_{\delta}(u)$,}$
where $B_{\delta}(u)$ is given in the $\|\cdot\|_{E}$-norm. We then infer that, for $\varepsilon\in(0,\varepsilon_{0})$,
$c_{\lambda+\varepsilon}$ $\leq\sup_{F^{+}(u)}\mathcal{J}_{\lambda+\varepsilon}=\sup_{F^{+}(u)\cap B_{\delta}(u)}\mathcal{J}_{\lambda+\varepsilon}$
$\leq\sup_{w\in F^{+}(u)\cap B_{\delta}(u)}\Bigl(\mathcal{J}_{\lambda}(w)-\frac{\varepsilon}{2}\|w\|_{H}^{2}\Bigr)$
$\leq\sup_{F^{+}(u)}\mathcal{J}_{\lambda}-\frac{\varepsilon}{2}\inf_{B_{\delta}(u)}\|w\|_{H}^{2}\leq c_{\lambda}-\frac{\varepsilon}{2}\Bigl(\|u\|_{H}-\delta\Bigr)^{2}$

### v3

# Normalized solutions of Nehari-Pankov type to mass-supercritical indefinite variational problems

@@ Abstract.
We consider abstract nonlinear equations of the form $Au=\lambda u+I^{\prime}(u)$, where $A$ is a self-adjoint operator with compact resolvent on a Hilbert space $H$, $\lambda\in\mathbb{R}$ is a parameter, and $u\mapsto I^{\prime}(u)$ is a superlinear term of variational nature. In this abstract setting, we develop a new approach to detect prescribed norm solutions in $H$ which does not rely on any mass-subcriticality assumptions. We then consider various applications of this approach. First, we obtain, under general assumptions including the full mass-supercritical parameter regime, the existence of (infinitely many) solutions to a class of nonlinear Schrödinger equations on a compact graph $\mathcal{G}$ with prescribed arbitrarily large mass, thereby improving previous results which only cover small masses. Moreover, we derive a similar result for a biharmonic Schrödinger equation in the $2$-torus. For a larger class of second order and higher order equations in a bounded domain with Dirichlet boundary conditions, we also show the existence of multiple solutions with prescribed small mass.
The solutions we obtain are detected as ground states of Nehari-Pankov type for the associated $\lambda$-dependent action functional, where $\lambda$ varies in a spectral gap between sufficiently large eigenvalues of $A$. The key new observation in this abstract framework is the fact that the $H$-norms of these $\lambda$-dependent solution families form connected sets even though the solution families themselves may be disconnected. To estimate the size of these connected sets in specific settings, we use Weyl type estimates for the length of spectral gaps, variational characterizations of eigenvalues, bounds for associated eigenfunctions and a bound from analytic number theory.
@@ 1. Introduction
The assumption (f1) is a natural superlinearity condition. The condition (f2) can be seen as a mild form of approximate homogeneity. Since, by (f1), the function $F$ is increasing on $\mathopen{[}0,+\infty\mathclose{)}$ and decreasing on $(-\infty,0]$, one can always replace $\kappa_{0}$ by a bigger positive constant in (4).
@@
Very recently, in [15, Theorem 1.3], Jeanjean and Song proved a multiplicity result for compact graphs $\mathcal{G}$ and small masses with continuity-Kirchhoff conditions. More precisely, by combining flow invariance techniques with delicate estimates, they show that for every $n\in\mathbb{Z}^{>0}$ there exists $\mu_{n}>0$ with the property that, for every $\mu\in(0,\mu_{n})$, the NLS equation $u^{\prime\prime}=\lambda u+|u|^{p-2}u$ admits at least $n$ pairs $\pm u$ of sign changing solutions with fixed mass $\mu$. Using our new “frequency to masses” tool (Theorem 1.8 below), we are able to improve this multiplicity result to the existence of infinitely many solutions, for any mass. Moreover, we can extend this multiplicity result beyond the class of odd nonlinearities to a setting where, in particular, Lusternik-Schnirelman theory does not apply. Let us also point out that we can deal with the most general self-adjoint vertex conditions, also generalizing the previous results. In the case of continuity-Kirchhoff conditions, the first eigenfunction of $A$ is positive in $\mathcal{G}$ (see e.g. [17, Theorem 4.12]), so that we also obtain the existence of sign-changing solutions. Finally, as already pointed out, we do not need to make any case distinction depending on whether $p<6$ or $6<p$.
@@
We emphasize that this result is new even in the classical second order case $m=1$, and it does not require oddness of the nonlinearity $f$ in $t$. In the special case of homogeneous nonlinearities of the form $f(x,t)=|t|^{p-2}t$, Theorem 1.3 is the analogue of the recent multiplicity result of Jeanjean and Song in [15, Theorem 1.3] for the case of compact metric graphs. Our method is completely different from the approach in [15], which, according to remarks in [15, Introduction], also yields multiplicity results for NLS-equations on bounded domains in $\mathbb{R}^{N}$.
@@
(i) In the mass supercritical regime $p>6$ and for a given positive mass $\mu>0$, even the mere existence of one normalized solution of (8) is new. If the weight function $r$ is constant, then the symmetries of the torus can be used to produce multiple solutions with large mass via a reflection procedure (see [20, Theorem 1.13 and Remark 1.1] for related results in bounded domains). This is not possible when $r$ does not have any symmetries.
@@
with strict inequality if $v=0$ and $s\neq 0$, see [24, Lemma 38]. This is very useful in applications.
@@
We emphasize that the connectedness of the set $M_{n}$ stated in Theorem 1.8 does not follow from a standard continuity argument, since we cannot expect connectedness of the set of associated ground state solutions in this general variational framework. Instead, the connectedness of $M_{n}$ follows by an inspection of the special saddle point nature of the Nehari-Pankov sets $\mathcal{N}\!{\mathcal{P}}_{\lambda}$, while it is not a sole consequence of the continuity of ground state levels $c_{\lambda}$.
@@
Let $k$ be a positive integer, and let $J_{1},\dots,J_{k}$ be disjoint spectral gaps [^3] We call an interval $J\subset\mathbb{R}$ a spectral gap of AA if $J=(\lambda_{n},\lambda_{n+1})$ for some $n\in\mathbb{Z}^{>0}$. of $A$. Then, there exists $\mu_{k}>0$ such that, for all $\mu\in(0,\mu_{k})$, there exist $\Lambda_{1},\dotsc,\Lambda_{k}\in\mathbb{R}$ with $\Lambda_{i}\in J_{i}$ for $i=1,\dots,k$ and $u_{1},\dotsc,u_{k}\in E$ with
@@
In particular, we are able to prove Theorem 1.1 by exploiting the presence of large spectral gaps for Laplacians on compact metric graphs (a consequence of Weyl’s law, which holds for all self-adjoint realizations of the Laplacian on graphs as shown in [6]). Classical spectral theory then shows that for parameters $\lambda$ at the bottom of a large spectral gap, solutions of (3) have a large $L^{\infty}$-norm. We then conclude using an ODE argument providing a suitable $L^{\infty}$ to $L^{2}$” estimate.
@@
The paper is organized as follows. In Section 2, we further develop the abstract framework of the indefinite variational problem associated with (1) and prove Theorem 1.7. In Section 3, we then study the $\lambda$-dependence of the Nehari-Pankov solutions obtained in Theorem 1.7, and we complete the proof of Theorem 1.8. In the end of Section 3, we then provide useful estimates for the action of Nehari-Pankov solutions under additional assumptions on the functional $I$ and the operator $A$. In Section 4, we then specialize to the case $H=L^{2}(\Omega)$ on a compact metric measure space and derive explicit lower bounds on the largest mass attainable by a Nehari-Pankov ground state within a given spectral gap, which underlie the large-mass results in Theorems 1.1 and 1.4. Section 5 is devoted to the nonlinear Schrödinger equation on compact metric graphs and to the proof of Theorem 1.1. Finally, in Section 6, we study the equation (8) on the flat $2$-torus, and we complete the proof of Theorem 1.4.
@@ 2. Existence of Nehari-Pankov solutions for fixed $\lambda$
Next, arguing as in [24, Chapter 4], we show that $\mathcal{N}\!{\mathcal{P}}_{\lambda}$ contains all nonzero critical points of $\mathcal{J}_{\lambda}$. Let $u\in E\setminus\{0\}$ be a critical point of $\mathcal{J}_{\lambda}$. Then, according to (20),
@@
(21) $\sup_{v\in E_{n}}\mathcal{J}_{\lambda}(v)\leq\sup_{v\in E_{n}}\frac{q_{\lambda}(v)}{2}\leq 0.$
@@
$w_{k}:=\frac{u_{k}}{t_{k}}\to w\qquad\text{in $E$ as $k\to\infty$,}$
@@
$\inf_{\mathcal{N}\!{\mathcal{P}}_{\lambda}}\mathcal{J}_{\lambda}=\inf_{\tilde{u}\in E\setminus E_{n}}\sup_{F_{n}(\tilde{u})^{+}}\mathcal{J}_{\lambda}.$
@@
Without loss of generality, we may assume that $w,w_{k}\in E_{n}^{\perp}$, and that $\|w\|_{E}=1$, $\|w_{k}\|_{E}=1$ for all $k$. We write
@@
$\mathcal{J}_{\lambda}(su+\tilde{v})=\lim_{k\to\infty}\mathcal{J}_{\lambda}(su_{k}+\tilde{v})\leq\lim_{k\to\infty}\mathcal{J}_{\lambda}(u_{k})=\mathcal{J}_{\lambda}(u).$
Consequently, $u$ maximizes $\mathcal{J}_{\lambda}$ on $F_{n}(u)^{+}=F_{n}(w)^{+}$, which implies that
@@ 3.1. Norms of Nehari-Pankov solutions
We first show (33). Let $\delta>0$ and $u\in Q_{\lambda}$. We claim that there exists $\varepsilon_{0}>0$ and $\kappa>0$ with
(35) $\mathcal{J}_{\lambda-\varepsilon}(w)\leq c_{\lambda}-\kappa\qquad\text{for all $\varepsilon\in[0,\varepsilon_{0}]$ and $w\in F_{n}(u)^{+}\setminus B_{\delta}(u)$,}$
@@
$\lim_{k\to\infty}\mathcal{J}_{\lambda-\varepsilon_{k}}(w_{k})\geq c_{\lambda}$
for a sequence $\varepsilon_{k}\to 0^{+}$ and a sequence $(w_{k})_{k}$ in $F_{n}(u)^{+}\setminus B_{\delta}(u)$. By Lemma 2.1 and the positivity of $c_{\lambda}$, the sequence $(w_{k})_{k}$ is bounded in $E$. Hence, since $F_{n}(u)^{+}\subset F_{n}(u)$ and the space $F_{n}(u)$ is finite dimensional, we may pass to a subsequence such that $w_{k}\to w$ in $E$ for some $w\in\overline{F_{n}(u)^{+}}\setminus B_{\delta}(u)$, which implies that
$\mathcal{J}_{\lambda}(w)=\lim_{k\to\infty}\mathcal{J}_{\lambda-\varepsilon_{k}}(w_{k})\geq c_{\lambda}>0.$
Consequently, $w\in F_{n}(u)^{+}\setminus B_{\delta}(u)$ by (21), but then the uniqueness stated in Proposition 2.2 gives that $u=w\in F_{n}(u)^{+}\setminus B_{\delta}(u)$, a contradiction. Hence (35) is true.
By the monotonicity property stated in Remark 3.1 and Proposition 2.2, we now observe that, for $\varepsilon\in(0,\varepsilon_{0})$,
$c_{\lambda}$ $\leq c_{\lambda-\varepsilon}\leq\sup_{F_{n}(u)^{+}}\mathcal{J}_{\lambda-\varepsilon}=\sup_{F_{n}(u)^{+}\cap B_{\delta}(u)}\mathcal{J}_{\lambda-\varepsilon}$
$\leq c_{\lambda}+\frac{\varepsilon}{2}\sup_{B_{\delta}(u)}\|\cdot\|_{H}^{2}\leq c_{\lambda}+\frac{\varepsilon}{2}\Bigl(\|u\|_{H}+\delta\Bigr)^{2}$
@@
From (33) and the fact that the map $\lambda\mapsto c_{\lambda}$ is decreasing, it immediately follows that this map is left continuous. Next, to prove its right continuity, let $\lambda\in(\lambda_{n},\lambda_{n+1})$, and let $(\lambda^{k})^{k}$ be a sequence in $(\lambda,\lambda_{n+1})$ which converges to $\lambda$. Since the map $\lambda\mapsto c_{\lambda}$ is decreasing, it follows that
$c_{\lambda}\geq\limsup_{k\to\infty}c_{\lambda^{k}}.$
Let, moreover, $w_{k}\in E_{n}^{\perp}$ be chosen with $\|w_{k}\|_{E}=1$ and $\sup\limits_{F_{n}(w_{k})^{+}}\mathcal{J}_{\lambda^{k}}=c_{\lambda^{k}}$. From Lemma 2.4, it then follows that
$w_{k}\rightharpoonup w\in E\qquad\text{after passing to subsequences}$
@@
$c_{\lambda}\leq\sup_{F_{n}(w)^{+}}\mathcal{J}_{\lambda}\leq\liminf_{k\to\infty}\sup\limits_{F_{n}(w_{k})^{+}}\mathcal{J}_{\lambda^{k}}=\liminf_{k\to\infty}c_{\lambda^{k}}.$
This shows that $\lim\limits_{k\to\infty}c_{\lambda^{k}}=c_{\lambda}$, and we thus obtain the full continuity of the map $\lambda\mapsto c_{\lambda}$.
Finally, to show (34), we again let $u\in Q_{\lambda}$ and $\delta\in\mathopen{(}0,\|u\|_{H}\mathclose{)}$. The $\varepsilon=0$ case of (35), combined with the monotonicity of $\lambda\mapsto\mathcal{J}_{\lambda}(w)$, gives $\mathcal{J}_{\lambda+\varepsilon}(w)\leq c_{\lambda}-\kappa$ for all $\varepsilon\geq 0$ and $w\in F_{n}(u)^{+}\setminus B_{\delta}(u)$, where $B_{\delta}(u)$ is given in the $\|\cdot\|_{E}$-norm. Moreover, by the continuity of $\lambda\mapsto c_{\lambda}$ established above, there exists $\varepsilon_{0}>0$ with $c_{\lambda+\varepsilon}>c_{\lambda}-\kappa$ for $\varepsilon\in(0,\varepsilon_{0})$. Using Proposition 2.2, we then infer that, for $\varepsilon\in(0,\varepsilon_{0})$,
$c_{\lambda+\varepsilon}$ $\leq\sup_{F_{n}(u)^{+}}\mathcal{J}_{\lambda+\varepsilon}=\sup_{F_{n}(u)^{+}\cap B_{\delta}(u)}\mathcal{J}_{\lambda+\varepsilon}$
$\leq\sup_{w\in F_{n}(u)^{+}\cap B_{\delta}(u)}\Bigl(\mathcal{J}_{\lambda}(w)-\frac{\varepsilon}{2}\|w\|_{H}^{2}\Bigr)$
$\leq c_{\lambda}-\frac{\varepsilon}{2}\inf_{B_{\delta}(u)}\|w\|_{H}^{2}\leq c_{\lambda}-\frac{\varepsilon}{2}\Bigl(\|u\|_{H}-\delta\Bigr)^{2}$